MATH 131, section 1 – Calculus for Physical and Life Sciences Recognizing f, f′, f′′ – October 10, 2007

In each of the following plots, one of the graphs is y = f(x), and the other two are y = f′(x) and y = f′′(x) (the original function’s first and second derivative functions). Which are which?

Graph I

Graph a is  , graph b is  , graph c is  .

Graph II

Graph a is  , graph b is  , graph c is  .